OpenStudy (anonymous):

Solve by factoring: x2-6x-16=0

OpenStudy (anonymous):

x^2-6x-16=0 (x+2)(x-8) = 0 x+2 = 0 or x-8 = 0 x = -2 or x = 8

OpenStudy (theviper):

\[\LARGE{x^2-6x-16=0}\]\[\Large{x^2-8x+2x-16=0}\]\[\Large{x(x-8)+2(x-8)}\]\[\Large{(x-8)(x+2)\Huge{\color{red}{\checkmark}}}\]

OpenStudy (anonymous):

thanks @TheViper you really do help me out alot :)

OpenStudy (theviper):

u r welcome brother :)